# Handling Label Noise via Instance-Level Difficulty Modeling and Dynamic Optimization

### v1

@@ Appendix B Additional Implementation Details
Regarding the hyperparameters of the model. In our experiments, we primarily utilized pre-trained ResNet-50 and ViT-16/B models, both of which were obtained by calling the PyTorch timm library. For the pre-training hyperparameters of the model, we adhered to the settings used in prior work [8, 22], as detailed in Table 6. One-stage baselines follow the setting in stage 2.
@@
Table 6: Optimizer configurations for different models and stages.
Regarding the details of the datasets. For synthetic datasets, we added noise to the entire training set and used the test set to evaluate performance. The model quality was determined by the accuracy on the test set at the last epoch. CIFAR100 consists of 100 classes, with 50,000 training images and 10,000 test images, and we set the batch size to 128. Tiny-ImageNet contains 200 classes, with 100,000 training images and 10,000 test images, and we also set the batch size to 128. For real-world datasets, we only used the noisy datasets and did not utilize the clean subsets provided by the datasets. Clothing1M is a class-imbalanced dataset, and we randomly sampled class-balanced subsets each time, with a batch size of 64 and 1,000 iterations. CIFAR100N is identical to CIFAR100 except for its real-world noisy labels. For WebVision, we used the first 50 classes, which include 66000 training images and 2,500 test images, and we set the batch size to 128.
@@
Appendix C Sensitivity Analysis of Parameter Configurations in Two Advanced Methods
C.1 Analysis and Experiments of ELR
ELR algorithm incorporates two crucial hyperparameters: the regularization term weight $\lambda$ and the temporal ensembling momentum coefficient $\beta$. Table 7 demonstrates the parameter sensitivity of ELR, showing that the performance is significantly influenced by parameter choices and comparison with the result of DODO. The parameter settings recommended for training from scratch in the original paper, where the authors recommend $\beta=0.7,\lambda=3$ for symmetric noise on CIFAR100 and $\beta=0.9,\lambda=7$ for asymmetric noise on CIFAR-100. However, parameter settings undergo significant shifts in pretrained settings. Optimal performance is achieved when $\beta\in[0.3,0.5],\lambda\in[3,5]$, demonstrating that optimal hyperparameters vary across different models and datasets, necessitating careful parameter tuning. Our proposed method DODO effectively addresses this challenge through dynamically weighted optimization.
Hyperparameter regularization term weight $\lambda$ momentum coefficient $\beta$ DODO
Noise Dataset 0 1 3 5 7 10 0.1 0.3 0.5 0.7 0.9 0.99
CIFAR-100 Sym. 60% 41.1 41.7 73.9 75.9 46.4 1.0 12.2 75.9 77.1 73.9 51.3 42.9 80.9
Table 7: Hyperparameter sensitivity of ELR, utilizing pre-trained ResNet50 on the CIFAR100 dataset with three noise settings. The default values are set as $\lambda=3$ and $\beta=0.7$. We consider the value ranges of $\lambda\in\{0,1,3,5,7,10\}$ and $\beta\in\{0.1,0.3,0.5,0.7,0.9,0.99\}$. following the discussion in original paper [10]
Hyperparameter $\mathcal{L_{U}}$ term weight $\lambda_{u}$ GMM threshold $\tau$ Mixup parameter $\alpha$ DODO
Noise Dataset 0 25 50 150 0.1 0.3 0.5 0.7 0.9 0.5 1 2 4
CIFAR-100 Sym. 60% 80.7 79.3 79.9 80.8 81.1 80.5 79.6 67.4 29.8 80.9 79.7 79.2 78.0 80.9
Table 8: Hyperparameter sensitivity of DivideMix, utilizing pre-trained ResNet50 on the CIFAR100 dataset with three noise settings. The default values are set as $\lambda_{u}=50$, $\tau=0.5$ and $\alpha=4$. We consider the value ranges of $\lambda_{u}\in\{0,25,50,150\}$, $\tau\in\{0.1,0.3,0.5,0.7,0.9\}$ and $\alpha\in\{0.5,1,2,4\}$ following the discussion in original paper [14].
C.2 Analysis and Experiments of DivideMix
DivideMix algorithm incorporates three crucial hyperparameters: the unsupervised loss term weight $\lambda_{u}$, GMM threshold $\tau$ for subset partitioning and beta distribution parameter $\alpha$ of Mixup [50]. Table 8 demonstrates the parameter sensitivity of DivideMix, showing that the performance is significantly influenced by parameter choices and comparison with DODO’s result. The parameter settings recommended for training from scratch in the original paper, where the authors recommend $\lambda_{u}=25,\tau\in[0.5,0.6],\alpha=4$ for symmetric noise on CIFAR100 and $\lambda_{u}=0,\tau=0.5,\alpha=4$ for asymmetric noise on CIFAR-100. However, parameter settings undergo significant shifts in pretrained settings. Optimal performance is achieved when $\lambda_{u}\in[50,150],\tau\in[0.1,0.5]$ shows that optimal hyperparameters vary between different models and data sets, which requires careful parameter tuning. Our proposed method DODO effectively addresses this challenge through dynamically weighted optimization.

### v2

@@ Appendix B Additional implementation details
Regarding the hyperparameters of the model. In our experiments, we primarily utilized pre-trained ResNet-50, ViT-16/B and ConvNeXt-B models, both of which were obtained by calling the PyTorch timm library. For the pre-training hyperparameters of the model, we adhered to the settings used in prior work [12, 13], as detailed in Table 8. One-stage baselines follow the setting in stage 2.
Table 8: Optimizer configurations for different models and stages.
@@
Regarding the details of the datasets. The model quality was determined by the accuracy on the test set at the last epoch. For synthetic datasets, we added noise to the entire training set and used the test set to evaluate performance. CIFAR100 consists of 100 classes, with 50,000 training images and 10,000 test images, and we set the batch size to 128. Tiny-ImageNet contains 200 classes, with 100,000 training images and 10,000 test images, and we also set the batch size to 128.
For real-world datasets, we only used the noisy datasets and did not utilize the clean subsets provided by the datasets. Clothing1M is a class-imbalanced dataset, and we sample class-balanced subsets each time, with a batch size of 64 and 1,000 iterations. Clothing1M-F means using fixed subset during training, used in [13, 12]. Clothing1M-R means sampling random subset per epoch, used in [16, 11]. In Clothing1M-R, we calculate wrong event for the subset every 100 iterations. CIFAR100N is identical to CIFAR100 except for its real-world noisy labels. For WebVision, we used the first 50 classes, which include 66000 training images and 2,500 test images, and we set the batch size to 128.
@@
Regarding the hyperpapameters of BMM. Since IDO relies on no priors, all experiments default that $B_{1}(1,2),B_{2}(2,1)$. The EM algorithm will update the parameters to the appropriate values.
Appendix C More analysis of wrong event
Figure 5: The baseline model is ResNet-18 trained on CIFAR-10 in 40% asymmetric noise for 100 epochs. We show the loss distributions (the first row) and wrong event distribution (the second row) during training. The four columns, (a) (b) (c) (d), represent the distributions at epoch 10, 50, 100 and the entire training phase. In (d), the heavy lines represent the median value and the shaded areas are the interquartile ranges, respectively. It is clear that wrong event can clearly separate noisy data and clean data, even if the model fits noise.
Figure 6: The baseline model is ResNet-18 trained on CIFAR-10 with 50% instance noise for 100 epochs. We show the loss distributions (the first row) and wrong event distribution (the second row) during training. The four columns, (a) (b) (c) (d), represent the distributions at epoch 10, 50, 100 and the entire training phase. In (d), the heavy lines represent the median value and the shaded areas are the interquartile ranges, respectively. It is clear that wrong event can clearly separate noisy data and clean data, even if the model fits noise.
C.1 More empirical observations of wrong event
We show a comparison between wrong event and loss under symmetric noise (see Fig.1). We also conduct experiments under instance noise, and asymmetric noise to verify that wrong event outperforms the loss across all types of noise. Fig.5 and Fig.6 show a comparison between wrong event and loss under asymmetric noise and instance noise, respectively. The experiment results reveal that in more challenging noisy environments, the model tends to rapidly fit the loss of noisy samples, resulting in compromised accuracy and stability of the provided modeling information. In contrast, wrong event metric maintains its capability to deliver robust modeling information about sample cleanliness and difficulty.
C.2 Analysis on the variance between wrong event and loss
We theoretically prove that loss variance has no upper bound, but wrong event does, decreasing with training. For epoch $T$, loss $l_{i}^{T}$, $l_{i}^{T+1}\in(0,+\infty)$, wrong event $w_{i}^{T+1}\in[w_{i}^{T},w_{i}^{T}+1]\subseteq[0,T]$, the variation is that
$\frac{w_{i}^{T+1}-w_{i}^{T}}{w_{i}^{T}}\in[0,1/T]\subseteq\frac{l_{i}^{T+1}-l_{i}^{T}}{l_{i}^{T}}\in(-1,+\infty)$
which clearly shows that the variance of wrong event has a upper bound, decreasing over training and the variance of loss has no upper bound, more likely to become large when the model already fits the given label but then predicts another label. We assume that loss variance becomes larger over training, because after the loss converge to loss, i.e., model remembers the sample, forgetting is more likely to bring a mutation to loss.
In experiments, the values of loss and wrong are not in the same range. We normalize metric values and track the range of value changes during training $\delta_{i}(l_{i}^{t})=|l^{t}_{i}-l^{t-1}_{i}|,\delta_{i}(w_{i}^{t})=|w_{i}^{t}-w_{i}^{t-1}|$ and $\text{Max Change}=\mathop{\text{argmax}}\limits_{(x_{i},\bar{y_{i}})\in\mathcal{D}}(\delta_{i}),\text{Avg Change}=\frac{1}{N}\sum_{i=1}^{n}\delta_{i}$. The result is in Figure 7. Loss variance is near the size of the dataset, confirming that loss has no bound. Wrong event variance decreases over training. Besides, due to the mutation of loss, the mean change is larger than wrong event. Experiment results confirm the theoretical analysis regardless of pre-training or not.
(a) Variance over epochs on pre-training setting
(b) Variance over epochs on random initialized setting
Figure 7: Variance comparison between loss and wrong event (WE). We test random pre-trained ResNet-50 on CIFAR-100 Inst. 40% noise (left) and random initialized ResNet-18 on CIFAR-10 Inst. 40% noise (right).
C.3 More discussion about different metrics
Researchers have proposed some metrics trying to replace loss because loss is easy to fit noise, such as forgetting event [18], fluctuation event [50], first k-epoch learning [19]. All of them are based on forgetting. In all the following equations, the symbol $\hat{y_{i}}^{t}$ is used to denote the model prediction of sample $(x_{i},\bar{y_{i}})$ at epoch $t$ Forgetting event (FE) is defined as
$\text{forgetting event}_{i}=\sum_{t=1}^{T}(\bar{y_{i}}=\hat{y_{i}}^{t-1})\wedge(\bar{y_{i}}\neq\hat{y_{i}^{t}})$
which counts how many times a model forgets a prediction of given labels. Fluctuation event is defined as
$\text{fluctuation event}_{i}=(\bar{y_{i}}=\hat{y_{i}^{t_{1}}})\wedge(\bar{y_{i}}\neq\hat{y_{i}^{t_{2}}})$
where $t_{1}\leq t_{2}$. First k-epoch learning (fkl) is defined as
$\text{fkl}_{i}=\mathop{\text{argmin}}\limits_{t^{*}\in t}[(\bar{y_{i}}=\hat{y_{i}}^{t^{*}})\wedge\cdots\wedge(\bar{y_{i}}=\hat{y_{i}}^{t^{*}-k+1})=1]$
which records the epoch when a sample has been predicted to its given label for consecutive k epochs for the first time.
Figure 8: The AUC-ROC curves of loss and wrong event. The experiment is conducted with random initialized ResNet-18 on CIFAR-10 (left) and pre-trained ResNet-50 on CIFAR-100 (right) under Inst.40% noise. The curves show the dynamic selection curves over epochs. We visualize the ability of clean sample selection in early, midterm and later training stages. The curves show that wrong event has a larger AUC value than loss and show good performance during the whole training phase, which represents a better ability of clean sample selection.
We group all three metrics into forgetting-based metrics (FBM). All metrics extract the statistical information to assist noise modeling, but are still different. Notably, FBM needs time to remember the given label, because the model needs time to fit the given label, especially when the label is noisy while wrong event needn’t. This difference makes FBM spend pretty longer time accumulating statistics. Considering easy noisy data for example, in early stage, model learns correctly and ground-truth of easy noisy data is consistently predicted, so the model cannot remember noisy labels. Easy noisy data’s wrong event would increase rapidly, but FBM would be constant value. Thus, FBM cannot distinguish easy noisy data in early stage, needing more training time and thus the scalability is limited, while wrong event has good modeling ability during the training phase. Table 4 also shows that loss-based metrics perform well in early stage but decrease over training, and FBM such as FE and fkl perform poorly in early stage but increase over training, needing more training time. While wrong event perform well during the entire training phase, showing effectiveness and efficiency.
C.4 Selection ability between wrong event and loss
Table 4 shows that wrong event outperforms loss in test accuracy. To further compare their clean sample selection ability, we calculate F-score along with precision and recall, AUC along with accuracy and precision under pre-training and random initialized settings. Table 9 and Table 10 show the F-score, Table 11 and Table 12 show the AUC value and Figure 8 shows the AUC-ROC curves during training. All results indicate that wrong event is better than loss in selecting clean samples. Wrong event can clearly model the noise during the whole training phase, while loss only performs well in early stage. In conclusion, wrong event is a simple but effective metric to select clean samples, which can capture noise well even if the model has fitted the noise.
Table 9: The precision, recall, F-score of loss and wrong event. The experiment is conducted with random initialized ResNet18 on CIFAR10 under Inst.40% noise. We set three thresholds to calculate F-score.
Threshold 0.2 0.5 0.8
Epoch 10 30 60 10 30 60 10 30 60
Loss F-score 0.49 0.38 0.29 0.34 0.29 0.20 0.00 0.00 0.00
Precision 0.62 0.76 0.79 0.71 0.79 0.75 0.92 0.90 0.60
Recall 0.41 0.25 0.18 0.23 0.18 0.11 0.00 0.00 0.00
Wrong Event F-score 0.60 0.67 0.70 0.63 0.70 0.72 0.70 0.79 0.76
Precision 0.43 0.51 0.54 0.46 0.54 0.56 0.63 0.88 0.93
Recall 0.99 0.99 0.99 0.99 0.99 0.99 0.77 0.71 0.65
Table 10: The precision, recall, F-score of loss and wrong event. The experiment is conducted with pre-trained ResNet-50 on CIFAR100 under Inst.40% noise. We set three thresholds to calculate F-score.
Threshold 0.2 0.5 0.8
Epoch 5 20 30 5 20 30 5 20 30
Loss F-score 0.77 0.84 0.56 0.83 0.61 0.04 0.05 0.01 0.00
Precision 0.63 0.75 0.41 0.81 0.97 0.94 1.00 1.00 1.00
Recall 1.00 0.96 0.87 0.85 0.45 0.02 0.02 0.00 0.00
Wrong Event F-score 0.76 0.85 0.84 0.85 0.90 0.92 0.87 0.92 0.88
Precision 0.62 0.76 0.74 0.73 0.83 0.89 0.83 0.93 0.97
Recall 1.00 0.97 0.99 0.99 0.98 0.96 0.93 0.92 0.81
Table 11: AUC value of wrong event and loss with random initialized ResNet-18 under different noise types and levels on CIFAR-10 (left) and CIFAR-100 (right). Higher AUC value indicates better clean sample selection ability.
Noise Sym. 60% Asym. 40% Inst. 40%
Epoch 20 60 100 20 60 100 20 60 100
Loss 0.95 0.80 0.63 0.80 0.64 0.58 0.84 0.65 0.59
Wrong Event 0.96 0.98 0.97 0.94 0.94 0.92 0.94 0.95 0.94
Noise Sym. 60% Asym. 40% Inst. 40%
Epoch 20 60 100 20 60 100 20 60 100
Loss 0.85 0.69 0.60 0.62 0.56 0.55 0.79 0.61 0.59
Wrong Event 0.87 0.92 0.91 0.72 0.75 0.74 0.86 0.88 0.88
Table 12: AUC value of wrong event and loss with pre-trained ResNet-50 under different noise types and levels on CIFAR-100. Higher AUC value indicates better clean sample selection ability.
Noise Sym. 60% Asym. 40% Inst. 40%
Epoch 5 15 25 5 15 25 5 15 25
Loss 0.96 0.88 0.74 0.66 0.63 0.59 0.84 0.75 0.67
Wrong Event 0.97 0.98 0.98 0.76 0.80 0.79 0.93 0.95 0.94
Appendix D Training time analysis
We analyze the training time of IDO to understand its efficiency. Results in Figure 4 show that IDO is faster than all baselines. For IDO itself, the running time $\approx$ BMM fitting + forward propogation. In Table 13, we compare the total training time of IDO on CIFAR-100 Inst.40% noise with several state-of-the-art methods which leverage probabilistic model, using a single Nvidia A100 80GB GPU.
IDO is faster than M-correction and DivideMix which involve multiple forward propagation and fitting iterations. In Table 13, we also break down the computation time for each operation in IDO. Results show that BMM is efficient, only 4.6s per epoch, which is 3.8% of the total training time. Moreover, results show that wrong event-based BMM in IDO is faster than loss-based BMM/GMM in M-correction and DivideMix. The reason is that the robustness of wrong event helps to reduce the fitting iteration rounds and require no reinitialization to save time. While loss often varies significantly in successive epochs, needing more fitting rounds and reinitialization in each epoch. In the open-sourced code, IDO only iterates few rounds based on the last BMM (1 iteration per round, no reinitialization), while M-correction and DivideMix need many rounds and reinitialization (10 GMM iterations in DivideMix, 20 BMM iterations in M-correction) to handle loss mutation.
Table 13: Comparison of methods using probabilistic model on CIFAR-100 with Inst. 40% Noise
Method Accuracy(%) Per Epoch Time(s) Per Epoch Fitting Time(s)
M-Correction(BMM) 78.9 399 14.7
DivideMix(GMM) 81.3 457 19.5
IDO(BMM) 83.7 121 4.6
Appendix E More exploratory experiments
In this section, we conduct more exploratory experiments, hoping to offer more insight and inspiration to the readers and community.
E.1 Base model vs. initial model
After Stage 1, we already have information about clean, noisy and hard samples. It is interesting to think that it is better to train the model from the base model or directly from the initial model. Since the base model has already been trained on corrupted samples, retraining from the initial model might lead to better performance.
Table 14: Comparison of initial model and base model on CIFAR-100 with different noise types
Start Model Sym. 60% Asym. 40% Inst. 40%
Initial Model 80.1 77.1 83.3
Base Model 81.5 78.0 83.7
We run experiments to verify this. Results are in Table 14. Overall, base model yields better results. Its advantage is high when handling Sym. noise as symmetric noise is harder to fit and easy to detect. The difference is smaller when handling Asym./Inst. noise as the model fits these noise more easily, decreasing the performance. Within the relatively limited training time, using the base model can more quickly achieve better performance.
E.2 Wrong event vs. loss accumulation
As the single epoch loss is easy to fit noise, it is interesting to think whether accumulated loss can prevent fitting noise. We have conducted new experiments to compare accumulated loss and wrong event. We get two observations from the results in Table 15.
Table 15: Comparison of wrong event and accumulated loss on CIFAR-100 with various noise types
Noise Sym. 60% Asym. 40% Inst. 40%
Single Loss + BMM 77.1 71.9 72.6
Accumulated Loss + GMM 79.2 74.5 77.9
Accumulated Loss + BMM 80.1 75.9 79.1
Wrong event + BMM 81.2 78.3 83.4
1. Accumulated loss performs better than single loss because accumulation prevents fitting noise, but the variance increases linearly when summing i.i.d. data. Besides, accidental sample forgetting can have a significant impact on the accumulated loss, increasing the variance. Thus, the distribution of loss sum is flatter than that of wrong event, which blurs the classification boundaries, making more data misclassified which drops the performance.
2. BMM performs better than GMM. In [16], the authors point out that GMM does not fit well on Asym. noise. Therefore, BMM, which shows stronger fitting ability, performs better.
E.3 Synthetic dataset vs. real world dataset
We find that IDO and all baselines achieve a smaller improvement on real-world datasets compared to synthetic datasets as shown in Table 1 and Table 2, which is worth further discussion.
We first analyze the difference between real-world noise and synthetic noise. [12, 51] pointed out that real world noise usually lies closer to the decision boundary than synthetic noise, which is harder to detect. In other words, it is easier to fit as boundary cases, slightly changing the decision boundary and thus results in less degradation of model performance than synthetic noise which will significantly change the decision boundary after fitted. As the degradation become smaller, the improvement space from w/ label noise to w/o label noise for algorithms is smaller.
E.4 Strong algorithm collapse in CIFAR-10
The experiment result for CIFAR-10 is in Table 16. IDO still outperforms baselines. It is surprising to see a performance drop outside of UNICON, a competitive algorithm in other datasets, showing that the variation of the hyperparameters across datasets and models easily decrease the model performance. This indicates high tuning cost and performance drop for hyperparameter-based methods in real world scenarios.
Table 16: Comparison with state-of-the-art LNL algorithms in test accuracy (%) on CIFAR-10.
Noise Architecture Sym. 20% Sym. 40% Sym. 60% Asym. 40% Inst. 40%
Standard ResNet-50 93.2% 92.3% 88.2% 91.1% 90.9%
UNICON ResNet-50 94.8% 93.2% 92.5% 93.5% 93.9%
ELR ResNet-50 96.5% 95.8% 95.1% 95.0% 94.8%
DeFT CLIP-ResNet-50 96.9% 96.6% 95.7% 93.8% 95.1%
DivideMix ResNet-50 97.1% 96.8% 96.3% 93.1% 96.0%
DISC ResNet-50 96.8% 96.5% 95.5% 95.2% 96.5%
[…]
Appendix F Limitations and broader impact
F.1 Limitations
Wrong event becomes suboptimal when the noise ratio becomes extremely high, i.e., more than 90%, because the distribution of wrong event will become extreme imbalance and the model struggles to converge. However, we do not suppose this will restrict the practical application ability of IDO, as real-world noise rates are often between 8.0% and 38.5% [52], far from extreme scenarios. However, designing a lightweight and highly scalable framework that remains competitive under extreme noise ratio is an important challenge for the community.
F.2 Broader Impact
Our research on label-noise learning contributes to building trustworthy machine learning systems, particularly in scenarios where high-quality annotations are costly or subjective (e.g., medical imaging diagnosis, social media content moderation). Consequently, as this method develops in effectiveness and scalability, the requirement for large-scale human-annotated data might decrease, potentially contributing to a rise in unemployment among data annotation professionals.